# references.bib

### v1

@@ IV-C Results
For evaluating computational cost, we measure the mean reconstruction time per object and the peak memory allocation during inference on the same data as used in our reconstruction accuracy evaluation. All our experiments are run on a server node with four 2.1GHz CPU cores, 20GB of RAM, and a NVIDIA RTX A6000 GPU with 48GB of VRAM. We note that the VRAM capacity is roughly in line with what current NVIDIA Jetson devices offer for on-robot compute. However, our GPU is significantly faster overall and consequently, the reported compute times would likely be significantly higher if the models were deployed on a NVIDIA Jetson. Figure 11 depicts that most evaluated models take multiple, sometimes tens of seconds for reconstructing an individual object. SF3D [boss_sf3d_2024] and ZeroShape [huang_zeroshape_2024] stand out by producing reconstruction within roughly 0.5 and 1 second, respectively. Still, reconstruction time scales linearly in the number of objects. We also observe large differences in memory allocation during inference. While SF3D [boss_sf3d_2024], LGM [tang_lgm_2025], Michelangelo [zhao_michelangelo_2023], and ZeroShape [huang_zeroshape_2024] all consume less than 10GB of VRAM, InstantMesh [li_instant3d_2023] and One2345 [liu_one-2-3-45_2023] require roughly triple and double the memory, respectively.
Comparing reconstruction error (Figures 2 and 13) and computational requirements (Figure 11), we do not observe a clear-cut trend that suggests computationally more expensive models, be it from having more parameters or generating multiple viewpoints, automatically produce more accurate reconstructions.
Scene reconstruction models require substantially more time to reconstruct their inputs than single-object reconstruction models (Figure 6). This appears sensible, as these models reconstruct multiple objects and also come with additional computational overhead to align reconstructed meshes. At the same time, VRAM utilisation is not too dissimilar to the single-object reconstruction models.
@@ V Discussion and Avenues for Future Research
In this paper, we evaluated the applicability of single-view 3D reconstruction models for creating digital twin environments in robot manipulation tasks. Our systematic assessment against key desiderata revealed significant limitations in current approaches when applied to robotics scenarios. Despite impressive advancements in computer vision benchmarks, these models struggle with the practical constraints of robotics applications, including low-error recostructions of typical robotic inputs, producing physically plausible and collision-free reconstructions, and managing occlusions in complex scenes.
The gap between computer vision capabilities and robotics requirements suggests several promising directions for future research. First, reconstruction models specifically trained on robotics-relevant data in addition to the current large-scale synthetic datasets could better address the input-domain-specific challenges we identified. In recent years, publications push for ever higher-resolution synthetic 3D datasets [deitke_objaverse-xl_2023, wu_omniobject3d_2023]. However, robot practitioners often require models that are robust to noise due to sensor limitations, changing illumination conditions, lower resolution inputs, and partial views from non-canonical angles. Rather than focusing solely on increasing the resolution and visual fidelity of synthetic 3D assets, future datasets should incorporate more diverse capture conditions that reflect real-world robotics scenarios. This includes training with simulated sensor noise patterns, varied lighting conditions, realistic occlusions, and the lower-resolution imagery typical of on-robot cameras. Additionally, datasets should include more examples of objects viewed from challenging angles—such as top-down or bottom-up perspectives—which proved particularly problematic in our evaluation.
@@
We advise the robot practitioner against using single-object reconstruction models if objects tend to be physically close in the scene. Single-object reconstruction models by design do not consider spatial contexts and thus are not able to avoid collisions between reconstructed meshes. Scene reconstruction models, on the other hand, are able to leverage spatial context and thus should be preferred for reconstructing multiple objects in a scene. However, our evaluation shows that even scene reconstruction models struggle with producing physically plausible reconstructions. One reason for this certainly is that some scene-level reconstruction approaches [dogaru_generalizable_2024, yao_cast_2025, chen_physgen3d_2025] rely on single-object reconstruction models under the hood. MIDI [huang_midi_2024] is an exception but does not show reduced object collisions either. A reason for this could be that while MIDI [huang_midi_2024] is trained on a scene-level dataset, the training data only contains larger objects such as room-level furniture, and using models trained on these datasets on tabletop manipulation tasks is likely to result in poor performance. This is a major limitation as most robotic manipulation tasks are performed on smaller objects, such as those in the YCB-Video [xiang_posecnn_2018] dataset. This highlights a common issue of scene-level datasets that often times come in one of two flavours: either they are real-world datasets that do not contain true 3D annotations (only 2.5D) [silberman_indoor_2012, dai_scannet_2017, chang_matterport3d_2017], or they are synthetic datasets resulting in a large domain gap between training and deployment environments [chang_shapenet_2015, deitke_objaverse-xl_2023]. In our view, the ideal training dataset for robotics-oriented reconstruction would include both synthetic data for scale and diversity, and carefully curated real-world scene data with accurate ground-truth meshes to bridge the gap between training and deployment environments.
@@
Object stability is another major limitation of current reconstruction models. While DSO [li_dso_2025] specifically trains for object stability, the authors set the threshold for object stability to 20 which is too large for most robotic manipulation tasks. Our evaluation shows that even the best-performing model, Michelangelo [zhao_michelangelo_2023], only achieves a median object stability of about 65% on YCB-Video [xiang_posecnn_2018] and 40% on the Meta Aria Digital Twin [pan_aria_2023] dataset. This is not sufficient for physical simulation, where object stability is crucial for scene integrity. We recommend that future models should be trained with a focus on object stability with tighter thresholds that are more in line with the requirements of robotic manipulation tasks.
@@
More generally, we suggest that future models should incorporate depth information more explicitly in the reconstruction pipeline. Depth information is heavily commoditised in robotics due to the prevalence of RGB-D cameras [tadic_perspectives_2022] and consequently widely used in applications [wei_d3roma_2024, li_visual_2024, kent_leveraging_2020]. Depth information can resolve location/scale ambiguities and can give almost ground truth supervision signal about object shapes. But, as most low-cost RGB-D cameras rely on structured light projection, depth information is often incomplete or noisy at locations that are occluded from the depth infrared sensor but not the RGB sensor, at highly reflective surfaces, low-texture regions, or affected by external light sources. With depth completion being non-trivial [zhang_deep_2018, khan_comprehensive_2022], reconstruction models should be able to cope with incomplete depth information, low depth resolution, and depth misalignment. None of the evaluated models allow utilising 2.5D information from RGB-D cameras in their reconstruction pipeline by design. One could argue that some parts of their reconstruction pipelines could be adapted to incorporate real-world depth information. For example, One2345 [liu_one-2-3-45_2023] could be adapted such that the dense 3D volume on which the denoising process is run is initialised with depth information. Michelangelo [zhao_michelangelo_2023] uses point clouds during pre-training to learn a latent space that merges shape, text, and image information. Again, one could consider initialising this latent space with point cloud information from observation before running the reverse diffusion process. Gen3DSR [dogaru_generalizable_2024] appears most readily adaptable since it relies on a separate pre-trained model to predict a depth map which is then utilised to align objects with the observation. However, during exploratory experiments where we replaced the depth prediction model with the observed depth map from the RGB-D camera, we found that Gen3DSR is unable to handle missing depth values -- a common issue with RGB-D cameras. Depth information offers a way of making the reconstruction process less ambiguous by providing information about the 3D structure of the scene. With depth information being commoditised in robotics, its utilisation in 3D reconstruction is a promising avenue towards more accurate reconstructions at almost no additional cost.
@@ Acknowledgments
The authors thank Joe Watson, Jack Collins, Jun Yamada, and Alex Mitchell for fruitful discussions and valuable assistance in the preparation of this manuscript. This work is supported by the European Laboratory for Learning and Intelligent Systems and a UKRI/EPSRC Programme Grant [EP/V000748/1]. We would also like to thank SCAN and ARC for use of their GPU acceleration facilities. Frederik Nolte is supported by the AWS Lighthouse Scholarship. Bernhard Schölkopf is a member of the Tübingen AI Center.
[…]
[heading=bibintoc,title=References]
VI Biography Section
Frederik Nolte is an ELLIS PhD student currently pursuing a DPhil in Engineering Science in the Applied AI Lab at the Oxford Robotics Institute. Previously, he earned an M.Sc. in Artificial Intelligence from the University of Amsterdam, where he was an ELLIS Honours Student.
Bernhard Schölkopf (Senior Member, IEEE) received degrees in mathematics (University of London, U.K., 1992), physics (University of Tübingen, Germany, 1994), and a Ph.D. in computer science (Technische Universität Berlin, Germany, 1997). He has researched at AT&T Bell Labs, Microsoft Research Cambridge, and Amazon. He is Director at the MPI for Intelligent Systems and at the ELLIS Institute Tübingen, Germany, and professor at ETH Zürich. His scientific interests include machine learning, causal inference and their applications across the sciences. Schölkopf received the Royal Society Milner Award, the Leibniz Prize, the Körber Prize, and the ACM AAAI Allen Newell Award.
Ingmar Posner (Member, IEEE) leads the Applied Artificial Intelligence Lab (A2I) at Oxford University and served as a Founding Director of the Oxford Robotics Institute (ORI). His research aims to enable machines to robustly act and interact in the real world - for, with, and alongside humans. It includes seminal work on large-scale learning from demonstration, unsupervised learning of scene dynamics and 3D object detection. Currently, Ingmar’s research focusses on representation learning for real-to-sim applications and world models in robotics and beyond. He is the recipient of a number of best paper awards at recognised international venues in robotics and AI such as ICAPS, IROS and ISER. An ELLIS Fellow, in 2014 he co-founded Oxa, a multi-award winning provider of mobile autonomy software solutions. He currently serves as an Amazon Scholar.
Appendix A Selected Reconstruction Models
Object-Level Reconstruction Models
SF3D [boss_sf3d_2024]: Extends TripoSR [tochilkin_triposr_2024] by adding material and lighting estimation. Uses a transformer to predict a Triplane [chan_efficient_2022], which is meshed via differentiable Marching Tetrahedron [shen_deep_2021].
InstantMesh [li_instant3d_2023]: Uses a multi-view diffusion model to generate views, which are encoded into Triplane [chan_efficient_2022] features. These features are then converted into meshes.
One2345 [liu_one-2-3-45_2023]: Produces dense multi-view images via diffusion, then aggregates them into a discretized volume. This is converted into a signed distance function and finally meshed.
LGM [tang_lgm_2025]: Generates multi-view images with diffusion and processes them via U-Net [ronneberger_u-net_2015] into 3D Gaussians. These are rendered to train a NeRF [mildenhall_nerf_2022], which is used for mesh extraction.
Michelangelo [zhao_michelangelo_2023]: Combines image, shape, and text in a latent autoencoder, followed by diffusion-based shape sampling. A conditioned occupancy network then extracts meshes.
ZeroShape [huang_zeroshape_2024]: Estimates depth and camera intrinsics to project the image into 3D space. A conditional occupancy network processes this for mesh generation.
Real3D [jiang_real3d_2024]: Builds on TripoSR [tochilkin_triposr_2024] and incorporates real-world data using an unsupervised loss. This co-training reduces domain shift and improves performance on real data.
DSO [li_dso_2025]: Fine-tunes TRELLIS [xiang_structured_2025] with a physics-informed reward system for stable 3D shapes. TRELLIS encodes images into structured latents, decoded into SDFs for mesh extraction.
Scene-Level Reconstruction Models
MIDI [huang_midi_2024]: Segments all objects in the scene using an image segmentation model. Then denoises latent embeddings jointly across objects using cross-instance self-attention, improving spatial reasoning. This enhances the model’s understanding of object relationships in the scene.
Gen3DSR [dogaru_generalizable_2024]: Segments and inpaints occluded objects before applying single-object 3D reconstruction. Aligns meshes to the scene using RANSAC [fischler_and_random_1981] between object meshes and estimated depth. Background reconstruction is omitted for fair comparison in experiments.
PhysGen3D [chen_physgen3d_2025]: Segments, inpaints, and reconstructs objects with InstantMesh [xu_instantmesh_2024]. Aligns them using keypoint descriptors and a differentiable renderer for refined pose. Adds language-model-inferred physical properties but does not handle collisions or stability during reconstruction.
[huang_midi_2024][dogaru_generalizable_2024][chen_physgen3d_2025]No CollisionCollision
Figure 12: Ratio of number of objects in collision to number of objects not in collision, averaged over a dataset of 28 YCB-Video [xiang_posecnn_2018] scenes. Even though scene reconstruction models have access to information about the entire scene, mesh collisions in 3D reconstruction are common. PhysGen3D [chen_physgen3d_2025] is more successful at avoiding object collisions. However, this could be an artifact from the model tending to markedly overestimate relative object distances (see Figure 11), thus erroneously pushing objects apart and avoiding collisions.
[boss_sf3d_2024][xu_instantmesh_2024][liu_one-2-3-45_2023][tang_lgm_2025][zhao_michelangelo_2023][huang_zeroshape_2024][jiang_real3d_2024][li_dso_2025]0123415Chamfer Distance (cm)
Figure 13: Reconstruction error on the Meta Aria Digital Twin [pan_aria_2023] dataset as measured by the Chamfer distance (CD). CD is averaged across 10,000 sampled surface points from both target and reconstruction. Indicators for min, max, and median are shown. Reconstruction errors on this dataset are noticeably higher than those shown in Figure 2. Reconstructed surfaces tend to be 1cm distant from the closest target mesh surface.
[boss_sf3d_2024][xu_instantmesh_2024][liu_one-2-3-45_2023][tang_lgm_2025][zhao_michelangelo_2023][huang_zeroshape_2024][jiang_real3d_2024][li_dso_2025]010203040100160Number of MeshesStableUnstable
Figure 14: Total number of objects with physically stable poses within 5 from the ground truth scene pose versus total number of objects without such stable poses. Evaluated on the Aria Digital Twin [pan_aria_2023] dataset. Similar to the results on the YCB-Video [xiang_posecnn_2018] dataset, most reconstructed object meshes across all methods lack the physical stability required for accurate simulation. When incorporated into physics engines, these unstable geometries cause entire scene reconstructions to collapse rather than maintain their intended configurations.
01234Chamfer Distance (cm)[huang_midi_2024][dogaru_generalizable_2024][chen_physgen3d_2025]-20-15-10-505101520Relative Error Increase (%)UnoccludedOccludedRelative Error Increase (%)
Figure 15: Chamfer distances (CD) of unoccluded object parts versus occluded object parts for scene-level reconstruction models. CD is averaged across occluded/unoccluded surface points from an initial set of 10,000 uniformly sampled surface points. To avoid error propagation from incorrect relative pose estimation on the reconstruction models’ part, we individually align each reconstructed object with its target mesh at the corresponding scene pose using the masked ICP procedure described in Section IV-B. Indicators for min, max, and median are shown. Gen3DSR [dogaru_generalizable_2024] and PhysGen3D [chen_physgen3d_2025] rely on image in-painting to resolve occlusions, whereas MIDI [huang_midi_2024] uses a custom shared denoising process to condition objects on one another. Both approaches result in noticeable improvements over single-object reconstruction models without any such procedures (see Figure 9). We conclude that scene-level information can be highly beneficial for resolving object occlusions.
Appendix B Further Results
Refer to captionSceneRefer to captionGTRefer to caption[boss_sf3d_2024]Refer to caption[xu_instantmesh_2024]Refer to caption[liu_one-2-3-45_2023]Refer to caption[tang_lgm_2025]Refer to caption[zhao_michelangelo_2023]Refer to caption[huang_zeroshape_2024]Refer to caption[jiang_real3d_2024]Refer to caption[li_dso_2025]Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix B Further Results
Figure 16: Sample reconstructions on the YCB-Video [xiang_posecnn_2018] dataset (rows 1, 2, 3) and the Meta Aria Digital Twin [pan_aria_2023] dataset (rows 4, 5, 6) using different 3D reconstruction methods. Reconstruction target objects are contoured and coloured. The scene renderings are solely for context as the reconstruction models are only given the segmented target object on transparent background. We observe that models tend to perform better on simple object shapes from highly informative viewpoints. Non-standard objects and more extreme viewpoints result in markedly lower reconstruction accuracy (compare rows 1, 2). Reconstruction quality on the Aria dataset is arguably lower than on YCB-Video. Objects in the Aria dataset tend to occupy less space on the image plane, resulting in less visual information in the model inputs.
Refer to captionSceneRefer to captionGTRefer to caption[huang_midi_2024]Refer to caption[dogaru_generalizable_2024]Refer to caption[chen_physgen3d_2025]Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 17: YCB-Video [xiang_posecnn_2018] scene reconstructions using different 3D reconstruction methods that process multi-object scenes. The objects that are reconstructed are shown contoured and in colour. The scene renderings are solely for context as the reconstruction models are only given the segmented target objects on transparent background. Scene reconstruction models struggle not only with object reconstruction quality but also accurate object pose estimation, resulting in mesh collisions and levitating objects.

### v2

@@ IV-C Results
For evaluating computational cost, we measure the mean reconstruction time per object and the peak memory allocation during inference on the same data as used in our reconstruction accuracy evaluation. All our experiments are run on a server node with four 2.1GHz CPU cores, 20GB of RAM, and a NVIDIA RTX A6000 GPU with 48GB of VRAM. We note that the VRAM capacity is roughly in line with what current NVIDIA Jetson devices offer for on-robot compute. However, our GPU is significantly faster overall and consequently, the reported compute times would likely be significantly higher if the models were deployed on a NVIDIA Jetson. Fig. 12 depicts that most evaluated models take multiple, sometimes tens of seconds for reconstructing an individual object. SF3D [80] and ZeroShape [142] stand out by producing reconstruction within roughly 0.5 and 1 second, respectively. Still, reconstruction time scales linearly in the number of objects. We also observe large differences in memory allocation during inference. While SF3D [80], LGM [96], Michelangelo [28], and ZeroShape [142] all consume less than 10GB of VRAM, InstantMesh [123] and One2345 [26] require roughly triple and double the memory, respectively.
Comparing reconstruction error (Figs. 3 and 14) and computational requirements (Fig. 12), we do not observe a clear-cut trend that suggests computationally more expensive models, be it from having more parameters or generating multiple viewpoints, automatically produce more accurate reconstructions.
Scene reconstruction models require substantially more time to reconstruct their inputs than single-object reconstruction models (Fig. 7). This appears sensible, as these models reconstruct multiple objects and also come with additional computational overhead to align reconstructed meshes. At the same time, VRAM utilisation is not too dissimilar to the single-object reconstruction models.
@@ V Discussion and Avenues for Future Research
We evaluated single-view 3D reconstruction for creating digital twin environments for robot manipulation tasks, revealing significant limitations despite computer vision advances. Despite impressive advancements in computer vision benchmarks, these models struggle with the practical constraints of robotics applications, including low-error reconstructions of typical robotic inputs, producing physically plausible and collision-free reconstructions, and managing occlusions in complex scenes.
The gap between computer vision capabilities and robotics requirements suggests several promising directions for future research. First, reconstruction models specifically trained on robotics-relevant data in addition to the current large-scale synthetic datasets could better address the input-domain-specific challenges we identified. In recent years, publications push for ever higher-resolution synthetic 3D datasets [48, 47]. However, robot practitioners often require models that are robust to noise due to sensor limitations, changing illumination conditions, lower resolution inputs, and partial views from non-canonical angles. Rather than focusing solely on increasing the resolution and visual fidelity of synthetic 3D assets, future datasets should incorporate more diverse capture conditions that reflect real-world robotics scenarios. This includes training with simulated sensor noise patterns, varied lighting conditions, realistic occlusions, and the lower-resolution imagery typical of on-robot cameras. Additionally, datasets should include more examples of objects viewed from challenging angles—such as top-down or bottom-up perspectives—which proved particularly problematic in our evaluation.
@@
We advise the robot practitioner against using single-object reconstruction models if objects tend to be physically close in the scene. Single-object reconstruction models by design do not consider spatial contexts and thus are not able to avoid collisions between reconstructed meshes. Scene reconstruction models, on the other hand, are able to leverage spatial context and thus should be preferred for reconstructing multiple objects in a scene. However, our evaluation shows that even scene reconstruction models struggle with producing physically plausible reconstructions. One reason for this certainly is that some scene-level reconstruction approaches [153, 36, 196] rely on single-object reconstruction models under the hood. MIDI [154] is an exception but does not show reduced object collisions either. A reason for this could be that while MIDI [154] is trained on a scene-level dataset, the training data only contains larger objects such as room-level furniture, and using models trained on these datasets on tabletop manipulation tasks is likely to result in poor performance. This is a major limitation as most robotic manipulation tasks are performed on smaller objects, such as those in the YCB-Video [38] dataset. This highlights a common issue of scene-level datasets that often times come in one of two flavours: either they are real-world datasets that do not contain true 3D annotations (only 2.5D) [212, 141, 172], or they are synthetic datasets resulting in a large domain gap between training and deployment environments [45, 48]. In our view, the ideal training dataset for robotics-oriented reconstruction would include both synthetic data for scale and diversity, and carefully curated real-world scene data with accurate ground-truth meshes to bridge the gap between training and deployment environments.
@@
Object stability is another major limitation of current reconstruction models. While DSO [190] specifically trains for object stability, the authors set the threshold for object stability to 20 which is too large for most robotic manipulation tasks. Our evaluation shows that even the best-performing model, Michelangelo [28], only achieves a median object stability of about 65% on YCB-Video [38] and 40% on the Aria Digital Twin [39] dataset. This is not sufficient for physical simulation, where object stability is crucial for scene integrity. We recommend that future models should be trained with a focus on object stability with tighter thresholds that are more in line with the requirements of robotic manipulation tasks.
@@
More generally, we suggest that future models should incorporate depth information more explicitly in the reconstruction pipeline. Depth information is heavily commoditised in robotics due to the prevalence of RGB-D cameras [67] and consequently widely used in applications [213, 214, 215]. Depth information can resolve location/scale ambiguities and can give almost ground truth supervision signal about object shapes. But, as most low-cost RGB-D cameras rely on structured light projection, depth information is often incomplete or noisy at locations that are occluded from the depth infrared sensor but not the RGB sensor, at highly reflective surfaces, low-texture regions, or affected by external light sources. With depth completion being non-trivial [216, 217], reconstruction models should be able to cope with incomplete depth information, low depth resolution, and depth misalignment. None of the evaluated models allow utilising 2.5D information from RGB-D cameras in their reconstruction pipeline by design. One could argue that some parts of their reconstruction pipelines could be adapted to incorporate real-world depth information. For example, One2345 [26] could be adapted such that the dense 3D volume on which the denoising process is run is initialised with depth information. Michelangelo [28] uses point clouds during pre-training to learn a latent space that merges shape, text, and image information. Again, one could consider initialising this latent space with point cloud information from observation before running the reverse diffusion process. Gen3DSR [153] appears most readily adaptable since it relies on a separate pre-trained model to predict a depth map which is then utilised to align objects with the observation. However, during exploratory experiments where we replaced the depth prediction model with the observed depth map from the RGB-D camera, we found that Gen3DSR is unable to handle missing depth values - a common issue with RGB-D cameras. Depth information offers a way of making the reconstruction process less ambiguous by providing information about the 3D structure of the scene. With depth information being commoditised in robotics, its utilisation in 3D reconstruction is a promising avenue towards more accurate reconstructions at almost no additional cost.
@@ Acknowledgments
The authors thank Joe Watson, Jack Collins, Jun Yamada, and Alex Mitchell for fruitful discussions and valuable assistance in the preparation of this manuscript. This work is supported by the European Laboratory for Learning and Intelligent Systems and a UKRI/EPSRC Programme Grant [EP/V000748/1]. We would also like to thank SCAN and ARC for use of their GPU acceleration facilities. Frederik Nolte is supported by the AWS Lighthouse Scholarship. Andreas Geiger and Bernhard Schölkopf are members of the Tübingen AI Center.
Appendix A Further Results
Refer to captionSceneRefer to captionGTRefer to caption[80]Refer to caption[27]Refer to caption[26]Refer to caption[96]Refer to caption[28]Refer to caption[142]Refer to caption[71]Refer to caption[190]Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix A Further Results
Figure 17: Sample reconstructions on the YCB-Video [38] dataset (rows 1, 2, 3) and the Aria Digital Twin [39] dataset (rows 4, 5, 6) using different 3D reconstruction methods. Reconstruction target objects are contoured and coloured. The scene renderings are solely for context as the reconstruction models are only given the segmented target object on transparent background. We observe that models tend to perform better on simple object shapes from highly informative viewpoints. Non-standard objects and more extreme viewpoints result in markedly lower reconstruction accuracy (compare rows 1, 2). Reconstruction quality on the Aria dataset is arguably lower than on YCB-Video. Objects in the Aria dataset tend to occupy less space on the image plane, resulting in less visual information in the model inputs.
Refer to captionSceneRefer to captionGTRefer to caption[154]Refer to caption[153]Refer to caption[196]Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 18: YCB-Video [38] scene reconstructions using different 3D reconstruction methods that process multi-object scenes. The objects that are reconstructed are shown contoured and in colour. The scene renderings are solely for context as the reconstruction models are only given the segmented target objects on transparent background. Scene reconstruction models struggle not only with object reconstruction quality but also accurate object pose estimation, resulting in mesh collisions and levitating objects.